# Perfecting Periodic Trajectory Tracking: Model Predictive Control with a Periodic Observer ($\Pi$-MPC)

### v1

@@ I Introduction
    Despite the use of simple models, our method consistently yields minimal tracking errors in both scenarios.
@@ II Problem Setup
$\begin{split}&x(t+1)=Ax(t)+Bu(t),\\ &y(t)=Cx(t),\end{split}$ (2)
with output $y\in\mathbb{R}^{n_{y}}$ and state $x\in\mathbb{R}^{n_{x}}$, allowing for $n_{x}$ to be different from $n$. We assume that $(A,B)$ is controllable, $(A,C)$ is observable, and $C$ has full row rank. The model is subject to the constraints
$x(t)\in\mathcal{X},\;u(t)\in\mathcal{U},\quad\forall t\in\mathbb{Z}_{\geq 0},$
where the sets $\mathcal{X}$ and $\mathcal{U}$ are assumed to be compact.
@@ III Periodic Disturbance Observer Design
In this section, we introduce a simple linear observer (7) to estimate periodic disturbances that account for the model mismatch. In particular, we first present an augmented model (4), discuss its observability (Prop. 1), and end by characterizing the observer’s convergence properties (Prop. 2).
@@ IV Periodic Model Predictive Control ($\Pi$-MPC)
We now present an MPC scheme that leverages the observer (7) to asymptotically achieve zero tracking error for periodic references. The observer provides estimates $\mathbf{\hat{d}}$ that are used to compute targets $\mathbf{\bar{x}}$, $\mathbf{\bar{u}}$ (11) for the state and input, respectively. The MPC problem (13) is formulated to minimize the deviation from these targets and ensure that the estimated controlled variable converges to the reference.
@@ IV-B $\Pi$-MPC formulation
where $Q\succeq 0$, $R\succ 0$, $(A,Q)$ is detectable, the terminal cost $P$ is chosen using a linear quadratic regulator (LQR), and we assume $L<N$ for simplicity. At each time $t$, the observer provides the state estimate $\hat{x}(t)$ and the estimates $\hat{d}_{k}(t)$ for the expected disturbance $k$ time steps ahead. We denote the optimal solution to (13) with a star ().
@@ VI-A Soft robot finite element simulation
    an MPC scheme using only the nominal model
@@
    the same MPC equipped with a constant disturbance observer (Offset-free MPC [8, 9, 10, 11])
@@
We conduct simulations through the SOFA finite-element-based physics simulator [27]. The mesh used to represent the Diamond robot is available in the SoftRobots plugin [28]. The Diamond robot features four actuators, shown in red in Fig. 1, each pulling at an elbow. The robot’s physical parameters match those reported in [29], with Young’s modulus of $E=175$ $\mathrm{M}\mathrm{P}\mathrm{a}$, Poisson ratio of $\nu=0.45$, and Rayleigh damping parameters $\alpha=2.5$ and $\beta=0.01$, where the damping matrix is defined as $C_{\mathrm{damp.}}=\alpha\mathbf{M}+\beta\mathbf{K}$. The output measurement, $y^{\text{f}}$, is the $xyz$ position of the robot’s tip (see 1) at time $k$ and at time $k-1$. A continuous-time, 6-dimensional linear model is learned using the framework outlined in [29] by specifying a first-order polynomial basis.
To build the MPC, we use a time-discretization of the model of $dt=0.01$ $\mathrm{s}$ and an MPC prediction horizon of $L=3$. Since the learned model’s $A$ matrix has no eigenvalues in common with $S_{d}$, we choose a disturbance model with $\overline{B}=0$, $\overline{C}=I$, see Remark 1. Finally, the observer gains in (7) are calculated using a Kalman filter, with resulting magnitudes of closed-loop eigenvalues between $0.68$ and $0.98$.
The control task has the robot’s end effector tracing a figure-eight along the $xy$–plane, with freedom in the $z$-direction. The figure-eight trajectory has an amplitude of $35$ $\mathrm{m}\mathrm{m}$ and a frequency of $2$ $\mathrm{H}\mathrm{z}$, corresponding to a period length of $N=50$. We note that no random noise is added to the simulation.
Fig. 1 shows the superior tracking performance of our method compared to the baselines over ten periods. The average tracking error in the last period is $32.66$ $\mathrm{m}\mathrm{m}$ for standard MPC, $18.59$ $\mathrm{m}\mathrm{m}$ for Offset-free MPC, and $0.25$ $\mathrm{m}\mathrm{m}$ for the proposed $\Pi$-MPC. The standard MPC scheme exhibits poor closed-loop tracking performance and fails to track the desired high-frequency trajectory, as the learned linear model does not give accurate predictions of the full nonlinear FEM system. The offset-free disturbance observer improves performance but still exhibits large tracking errors as it tries to estimate a constant disturbance, despite the model mismatch being time-varying. Instead, our approach properly considers the disturbances at each point throughout the trajectory. As expected from the presented theory, $\Pi$-MPC ensures that the tracking error decays to zero asymptotically despite significant model discrepancy. In fact, after 50 periods, the error reduces below $1\times 10^{-2}$ $\mathrm{m}\mathrm{m}$.
@@ VI-B Miniature race car experiments
In the experiments, we compare a naïve MPC implementation, using only the model, with the proposed $\Pi$-MPC approach, which additionally uses the periodic disturbance observer. The experimental results are illustrated in Fig. 2. Both MPC formulations provide identical results in the first lap. After ten laps, the MPC baseline has an average error of $6.22$ $\mathrm{c}\mathrm{m}$, while $\Pi$-MPC has an average error of only $1.26$ $\mathrm{c}\mathrm{m}$, five times lower. Over the course of sixteen laps, the proposed formulation reduces the peak tracking error in a lap to $2.8$ $\mathrm{c}\mathrm{m}$ while the naïve MPC implementation continues to show peak errors over $14$ $\mathrm{c}\mathrm{m}$. Hence, as seen in Fig. 2, we reduce the tracking error by a factor of five after only a few laps. Although the theory suggests convergence to zero error, the presence of non-deterministic effects, such as noise and delays, may result in small fluctuations.
@@ VII Conclusion
  • easily implemented on top of an existing MPC scheme,

### v2

@@ I Introduction
    Despite the use of simple models, our method consistently achieves significantly lower tracking errors, reducing them by factors ranging from 4 to over 5000 compared to baseline MPC.
@@ II Problem Setup
$\begin{split}x(t+1)=Ax(t)+Bu(t),\;\;y(t)=Cx(t),\end{split}$ (2)
with output $y\in\mathbb{R}^{n_{y}}$ and state $x\in\mathbb{R}^{n_{x}}$, allowing for $n_{x}$ to be different from $n$. We assume that $(A,B)$ is controllable, $(A,C)$ is observable, and $C$ has full row rank. The model is subject to the constraints $x(t)\in\mathcal{X},\;u(t)\in\mathcal{U},\;\forall t\in\mathbb{Z}_{\geq 0},$ where the sets $\mathcal{X}$ and $\mathcal{U}$ are assumed to be compact.
@@ III Periodic Disturbance Observer Design
In this section, we introduce a simple linear observer (7) to estimate periodic disturbances. These estimated disturbances should compensate for the deterministic model mismatch. In particular, we first present an augmented model (4), discuss its observability (Prop. 1), and end by characterizing the observer’s convergence properties (Prop. 2).
@@ IV Periodic Model Predictive Control ($\Pi$-MPC)
We now present an MPC scheme that leverages the observer (7) to asymptotically achieve zero tracking error for periodic references. The observer provides estimates $\mathbf{\hat{d}}$ that are used to compute targets $\mathbf{\bar{x}}$, $\mathbf{\bar{u}}$ (11) for the state and input, respectively. The MPC problem (13) is then formulated to minimize deviations from these targets, ensuring the estimated controlled variable converges to the reference.
@@ IV-B $\Pi$-MPC formulation
where $Q\succeq 0$, $R\succ 0$, $(A,Q)$ is detectable, the terminal cost $P$ is chosen using a linear quadratic regulator (LQR), and we assume $L<N$ for simplicity [^4] Otherwise we repeat the periodic targets to fill the horizon length.. At each time $t$, the observer provides the state estimate $\hat{x}(t)$ and the estimates $\hat{d}_{k}(t)$ for the expected disturbance $k$ time steps ahead. We denote the optimal solution to (13) with a star ().
@@ VI-A Soft robot finite element simulation
    an MPC scheme using only the nominal model,
@@
    the same MPC equipped with a constant disturbance observer (Offset-free MPC, OF-MPC [8, 9, 10, 11]),
@@
TABLE I: Quantitative Evaluation of Tracking Errors in Soft-Robot Sim.
10th Period 50th Period
Avg. Max. Avg. Max.
MPC 30.5 $\mathrm{m}\mathrm{m}$ 46.8 $\mathrm{m}\mathrm{m}$ 30.0 $\mathrm{m}\mathrm{m}$ 47.4 $\mathrm{m}\mathrm{m}$
OF-MPC 18.6 $\mathrm{m}\mathrm{m}$ 28.7 $\mathrm{m}\mathrm{m}$ 18.6 $\mathrm{m}\mathrm{m}$ 28.7 $\mathrm{m}\mathrm{m}$
$\mathbf{\Pi}$-MPC 0.26 $\mathrm{m}\mathrm{m}$ 0.52 $\mathrm{m}\mathrm{m}$ 0.004 $\mathrm{m}\mathrm{m}$ 0.008 $\mathrm{m}\mathrm{m}$
We conduct simulations through the SOFA finite-element-based physics simulator [27]. The Diamond robot mesh is available in the SoftRobots plugin [28]. The robot features four actuators, shown in red in Fig. 1, each pulling at an elbow. The robot’s physical parameters match those reported in [29]. The output measurement, $y^{\text{f}}$, is the $xyz$ position of the robot’s tip at time $k$ and at time $k-1$. Following the framework outlined in [29], we regress a continuous-time linear model by collecting decaying trajectories and identifying a low-dimensional subspace with latent coordinates $x(t)\in\mathbb{R}^{6}$.
For MPC, we discretize the model with a time step of $dt=0.01$ $\mathrm{s}$. To achieve real-time control, we select a prediction horizon of $L=15$ steps. This configuration results in solve times of approximately 3 $\mathrm{m}\mathrm{s}$ across all MPC schemes. Since the learned model’s $A$ matrix has no eigenvalues in common with $S_{d}$, we choose a disturbance model with $\overline{B}=0$, $\overline{C}=I$, cf. Remark 1. Finally, the observer gains in (7) are calculated using a Kalman filter, with resulting magnitudes of closed-loop eigenvalues between $0.68$ and $0.98$.
The control task has the robot’s end effector tracing a figure-eight along the $xy$–plane, with freedom in the $z$-direction. The figure-eight trajectory has an amplitude of $35$ $\mathrm{m}\mathrm{m}$ and a frequency of $2$ $\mathrm{H}\mathrm{z}$, corresponding to a period of $N=50$. No random noise is added to the simulation.
Fig. 1 illustrates the superior tracking performance of our method relative to the baseline methods across ten periods, with quantitative results detailed in Table I. The standard MPC scheme exhibits poor closed-loop tracking performance and fails to track the desired high-frequency trajectory, as the learned linear model does not give accurate predictions of the full nonlinear FEM system. The offset-free disturbance observer improves performance but still exhibits large tracking errors as it tries to estimate a constant disturbance, despite the model mismatch being time-varying. Instead, our approach properly considers the disturbances at each point throughout the trajectory. As expected from the presented theory, $\Pi$-MPC ensures that the tracking error decays to zero asymptotically despite significant model discrepancy. In fact, after 50 periods, the peak error reduces below $1\times 10^{-2}$ $\mathrm{m}\mathrm{m}$.
@@ VI-B Miniature race car experiments
In the experiments, we compare a naïve MPC implementation, using only the model, with the proposed $\Pi$-MPC approach, which additionally uses the periodic disturbance observer. The experimental results are illustrated in Fig. 2. Quantitative results are summarized in Table II. Both MPC formulations provide identical results in the first lap. After ten laps, $\Pi$-MPC has an average error approximately 4 times lower than the MPC baseline. Over the course of sixteen laps, the naïve MPC implementation continues to show peak errors over $14$ $\mathrm{c}\mathrm{m}$, while the proposed formulation reduces the peak tracking error below $3$ $\mathrm{c}\mathrm{m}$. Although the theory suggests convergence to zero error, the presence of non-deterministic effects, such as noise and delays, may result in small fluctuations.
TABLE II: Quantitative Evaluation of Tracking Errors in Race Car Exp.
5th Lap 16th Lap
Avg. Max. Avg. Max.
MPC 6.19 $\mathrm{c}\mathrm{m}$ 14.21 $\mathrm{c}\mathrm{m}$ 6.15 $\mathrm{c}\mathrm{m}$ 14.23 $\mathrm{c}\mathrm{m}$
$\mathbf{\Pi}$-MPC 5.54 $\mathrm{c}\mathrm{m}$ 8.04 $\mathrm{c}\mathrm{m}$ 1.42 $\mathrm{c}\mathrm{m}$ 2.91 $\mathrm{c}\mathrm{m}$
@@ VII Conclusion
  • easily implemented on top of an existing MPC scheme with both linear or nonlinear models,